# Explainable AI for computational pathology identifies model limitations and tissue biomarkers

### v2

@@ Results
We hypothesized that the low specificity of the UNI-based ABMIL model may be attributed to metastasis-induced alterations in the surrounding tumor microenvironment. To investigate this, we used HIPPO-knowledge to remove increasingly larger regions surrounding the tumor annotation and quantified the effect on metastasis detection. As the extent of peritumoral tissue removal increased, the UNI-based model was consistently more likely to predict the absence of metastasis. Specificity increased from 0.73 at dilation of , to 0.78 at , to 0.80 at , to 0.86 at , and to 0.88 at . This was driven primarily by macrometastatic specimens, where specificity increased from 0.59 at dilation of to 0.68 at dilation of , to 0.73 at dilation of , to 0.82 at dilation of , to 0.86 at dilation of . Notably, other ABMIL models remained largely unaffected by peritumoral tissue removal, highlighting a unique characteristic of the UNI-based model (Supplementary Fig. 1). In summary, HIPPO enabled the quantitative exploration of peritumoral tissue on metastasis detection.
@@
While necessity assesses the importance of a feature or feature set, it does not inform whether the feature set is sufficient for model predictions. Metastasis detection models must be able to detect tumor regions no matter how small. Using the HIPPO-knowledge method, we tested the sufficiency of metastatic regions using two methods: removing all non-tumor patches and measuring model outputs and adding tumor regions to normal specimens and measuring model outputs.
First, we constructed counterfactual specimens (n=49) by removing all non-tumor tissue (i.e., removing patches that did not intersect with expert tumor annotations) and measuring model outputs. With only the tumor present, the true label for these images was “positive”, and the foundation models had the following sensitivity (true positive rate): UNI-based 0.98, REMEDIS-based 0.92, Phikon-based 0.98, CTransPath-based 0.96, RetCCL-based 0.82 (Fig. 2d). There is evidence to suggest that extratumoral tissue caused false negative predictions. Four of the five foundation models improved sensitivity when using only tumor tissue in micrometastases compared to the original positive samples, suggesting that extratumoral tissue drove false negative predictions. The sensitivity of CTransPath increased by 25%, Phikon by 4%, REMEDIS by 5%, and RetCCL by 100%. For UNI, however, using original WSIs resulted in a sensitivity of 1.0 on micrometastasis. However, when using only the tumor tissue, one false negative prediction suggested that the UNI-based model may use tissue outside of the metastatic region in its predictions. Critically, this demonstrated that the tumor was insufficient for a positive prediction in this specimen with the UNI-based model and that extratumoral tissue was solely driving the positive prediction. RetCCL had a true positive rate in macrometastases of 0.95 (21 predicted positive of 22 positive specimens). When using only tumor tissue, all macrometastases were detected successfully, demonstrating that tissue outside the metastatic region caused a false negative prediction.
@@
We also evaluated whether tumor was sufficient for metastasis detection by embedding tumor regions in normal specimens. We embedded all patches intersecting with tumor annotations into normal specimens, resulting in 3,920 positive counterfactual examples (80 normal slides 49 positive slides). Model outputs for these examples were recorded. The UNI-based model had a sensitivity of 0.98, REMEDIS-based 0.86, Phikon-based 0.95, CTransPath-based 0.90, and RetCCL-based 0.63. Positive counterfactuals made with micrometastases were less likely to be detected by most models (UNI-based achieved sensitivity of 0.96, REMEDIS-based 0.75, Phikon-based 0.91, CTransPath-based 0.93, and RetCCL-based 0.40), suggesting that smaller tumors in the context of normal tissue are insufficient for positive metastasis detection (Fig. 2e).
@@
The average treatment effect for each metastatic slide was calculated by averaging the model’s probability of metastasis across all negative samples. This informs which positive slides can drive positive predictions across individuals. 100% of macrometastases (n=22) led to true positives in UNI-based, REMEDIS-based, Phikon-based, and CTransPath-based models. In the RetCCL-based model, 90% (n=20) of macrometastases had an average true positive effect. Micrometastases (n=27) were less likely to induce positive predictions on average, with 96% (n=26) positive in UNI, 93% (n=25) in Phikon, 81% (n=22) in CTransPath, 74% (n=20) in REMEDIS, and 37% (n=10) in RetCCL.
@@
To evaluate the sensitivity of ABMIL models to detect metastasis based on the size of the metastasis in a specimen, we analyzed the metastasis-positive specimens from the CAMELYON16 test set. Our methodology involved initially removing all tile embeddings that intersected with expert tumor annotations, effectively rendering the slide negative for metastases. A region of tumor (shown in the right-hand side of Fig. 2a) was added to 80 normal specimens and 49 metastasis-removed positive specimens. When the single-patch tumor region was embedded in normal specimens, the REMEDIS-, Phikon-, and RetCCL-based ABMIL models detected 100% of counterfactuals as positive, highlighting their robustness to this small region of tumor. The UNI-based model, on the other hand, failed to detect 41% (n=33) of positive counterfactuals (n=80), and the CTransPath-based models failed to detect 35% (n=28) of positive counterfactuals. A similar trend was observed when the tumor region was embedded into the context of metastatic specimens (i.e., the positive specimen with metastasis removed). The REMEDIS-, Phikon-, and RetCCL-based models detected 100% of positive counterfactuals (n=49), whereas the UNI-based model missed 51% (n=25) and CTransPath-based missed 65% (n=32) of positive counterfactuals specimens (Fig. 2f). This result is surprising because the UNI-based model had perfect sensitivity in the original test set (Fig. 2b) as well as the highest sensitivity when larger tumors were embedded into normal tissue (Fig. 2e). This highlights that high classification performance on the held-out test set is insufficient to assess generalization to more nuanced downstream applications.
@@
We also sought to quantify the sensitivity of models to each tumor patch in positive specimens, which can shed light on whether tumor patches carry different levels of informativeness for machine learning classifiers. To accomplish this, all tumor patches intersecting with expert tumor annotations were removed. Then, we reintroduced tiles fully within the expert tumor annotation, one at a time, to the tumor-removed specimen and evaluated the model outputs. These model outputs were compared to those when all tumor was removed. While some tumor patches could drive a positive prediction on their own, many could not (Fig. 2g for the UNI-based model, and other models and specimens are shown in Supplementary Figs. 2-6.).
@@
To further quantify the effect of tumor size in metastasis detection, we added tumor patches into normal slides in a graded fashion and measured the sensitivity. All models exhibited a graded effect of tumor size, and UNI exhibited the highest sensitivity (Fig. 2h). Models tended to plateau in sensitivity at of tumor (16 patches) added. The RetCCL-based model showed the lowest sensitivity and the least sensitivity to smaller tumors.
@@
To identify the largest amount of tumor that would go undetected by a model, we also used a HIPPO search method, HIPPO-search-low-effect. We found that in some cases, regions up to could be added into a negative counterfactual while still maintaining a negative detection. Indeed, the tumor patches that were insufficient to drive large effect sizes were largely similar to the sufficient patches, though some insufficient patches contained adipose cells along with tumor epithelial cells (Supplementary Fig. 7). This shows that there exist regions within tumors that would go unseen by an ABMIL-based metastasis detection model. These biases should be explored further prior to the clinical use of metastasis detection models.
@@
Given the effect of peritumoral tissue on UNI-based model predictions, we also evaluated whether peritumoral tissue was sufficient for positive metastasis predictions. A halo of peritumoral tissue was extracted from metastasis-positive specimens (n=49) with a width of 64, 128, 256, or 1024 , beginning at either the edge of the expert tumor annotation or outside of the tumor annotation. The patches intersecting with the peritumoral halos were added to normal specimens (n=80), resulting in 3920 counterfactual examples (80 normal 49 positive specimens). Model predictions were averaged across normal specimens, resulting in the average treatment effect of the peritumoral region of each positive specimen. This was evaluated for the UNI-based and Phikon-based models.
@@
Halos of peritumoral tissue were sufficient to drive positive metastasis detection. Halos in width beginning from the edge of the expert tumor annotation from 20% (n=10) of positive specimens caused positive predictions in the UNI-based model when embedded in normal specimens, whereas 10% (n=5) caused positive predictions in the Phikon-based model. When starting the halo outside of the tumor annotation, there were 14% (n=7) positive predictions in the UNI-based model and 10% (n=5) for the Phikon-based model. Thinner halos of beginning outside of the tumor annotation also caused positive predictions: 10% (n=5) in the UNI-based model and 6% (n=3) in the Phikon-based model (Supplementary Fig. 8). These results demonstrate that the models learned an association between peritumoral tissue and the presence of metastasis despite the peritumoral tissue not containing any metastasis. This highlights an important bias that could not have been uncovered using attention alone. HIPPO enabled the quantitative assessment of peritumoral tissue on metastasis detection in multiple foundation models.
Adipose tissue can cause false negative metastasis detections
To investigate patient samples that consistently drove false negative predictions further, we visualized the attention maps of the ABMIL models to identify regions that it considers important. For the CTransPath-based model, we observed that attention was concentrated in adipose regions for specimen test_066 (Figs. 3a and 3b).
@@
Since attention maps only provide a qualitative visualization of regions in an image that the ABMIL models consider important, it is unclear to what extent adipose tissue directly affects model predictions. We address this with HIPPO-attention. The patches with adipose tissue were removed, and the effect of this perturbation on model outputs was quantified. We found that removing the adipose tissue rescued the true positive prediction in specimen test_066, suggesting that fat caused the false negative prediction (Figs. 3c and 3d). To test whether the adipose tissue from test_066 would cause a misclassification in other models and specimens, we added the adipose regions from that specimen into the 48 other positive specimens. We then recorded whether the addition of adipose tissue caused false negative predictions. We found that true positives were flipped to false negatives in 2, 2, 1, 1, and 5 specimens for UNI, REMIDIS, Phikon, CTransPath, and RetCLL, respectively (Figs. 3e and 3f). This highlights how HIPPO can elucidate biases that cause misclassification. While attention alone could not inform that adipose was the cause of misclassification, it was useful to formulate a hypothesis that we can test with HIPPO-attention. This demonstrates how HIPPO can complement attention-based interpretability analysis to quantitatively test hypotheses of putatively important tissue regions.
@@
Identifying spurious correlations in deep learning models for medical imaging is crucial to ensure reliable and clinically relevant results. To test HIPPO’s ability to identify spurious correlations, we conducted an experiment where we deliberately introduced an artificial bias into the CAMELYON16 dataset (Supplementary Figs. 9a and 9b). Specifically, blue squares were added to all negative images. This mimics the plausible scenario in which a pathologist marks certain slides with a blue marker. However, in doing so, it introduces a strong spurious correlation with labels. We hypothesized that the models would learn that slides were negative if a blue region was present and that slides lacking this blue region are positive (as blue regions are easier to identify compared to more variable tumor regions).
@@
An ABMIL model was trained on the modified training data using UNI embeddings. The model achieved a balanced accuracy of 1.0 on the test set, suggesting the spurious correlations created a trivial prediction task. By performing standard model interpretation using attention, we found that metastatic regions were considered highly important (Supplementary Fig. 9c). However, removing these regions using HIPPO did not alter the model predictions, demonstrating that tumor regions were not important for model predictions despite a strong attention assignment. This highlights an important weakness of attention: the disconnect between attended regions and model predictions.
@@
Knowing that the metastatic regions did not affect model outputs, we used the search algorithm HIPPO-search-high-effect to identify the regions that maximally drove positive tumor predictions in both models using one positive specimen, test_002. Given that the model trained with spurious correlations uses the lack of a blue square as a cue for positive specimens, we expected that no individual patches would drive the positive metastasis output and that tumor regions would not have a high effect on the prediction. Indeed, effect sizes were small and evenly distributed across the WSI (minimum , maximum , mean , and median ), indicating that no single region contributed strongly to the model prediction (Supplementary Fig. 9d). By contrast, applying this search algorithm to the model trained on the original CAMELYON16 dataset, we found that patch effect sizes were higher (minimum , maximum , mean , and median ), and high effect patches were within expert tumor annotations (Supplementary Fig. 9e). By tying interpretation analysis directly to predictions, HIPPO-based interpretations may provide more reliable explanations of model predictions.
Shortcut learning is an important bias that must be identified and addressed in deep learning on medical images. In this case, model performance and attention were insufficient to diagnose the shortcut learning. Observational analysis based on attention maps could easily mislead an observer to believe that tumor regions drive model predictions. Quantifying effect sizes of tumor regions using HIPPO addressed these limitations and diagnosed the shortcut learning.
@@
Having demonstrated HIPPO’s effectiveness in metastasis detection, where the regions of interest are well-defined and were previously annotated by expert pathologists, we extended our investigation to the more complex domain of cancer prognosis. Unlike the clear delineation of tumor regions in metastasis detection, prognostic factors in WSIs are multifaceted and less clearly defined. We applied HIPPO to prognostic models that generate risk scores from WSIs, aiming to identify the tissue regions driving these predictions. Our experiments with HIPPO yielded two key insights. First, HIPPO’s search algorithms demonstrated superior ability in identifying tissue patches that consistently and significantly influence risk predictions compared to conventional attention-based methods. While attention mechanisms yielded mixed effects — potentially identifying regions that counterintuitively drive lower risk in otherwise high-risk specimens — HIPPO provided a more consistent, reliable, and quantitative assessment of the regions that drive risk. Second, HIPPO’s unique features enable virtual experiments to measure the effects of targeted tissue interventions on prognostic outcomes through the lens of the ABMIL model. HIPPO’s potential to accelerate the discovery and validation of prognostic tissue biomarkers is an exciting development in cancer research, potentially bridging the gap between computational predictions and clinical actionability.
@@
We trained prognostic ABMIL models using the PORPOISE framework [50], a computational tool designed for predicting survival outcomes from histopathology images, to predict overall survival from WSIs in breast cancer (TCGA-BRCA) and cutaneous melanoma (TCGA-SKCM) (Supplementary Fig. 10). The same training and validation splits were used as in the original publication. Non-overlapping patches from WSIs were embedded using the UNI model [20] (in the original PORPOISE publication, a truncated ResNet50 [88] was used). Low and high risk were defined as the first and fourth quartiles of risk scores. High attention regions were defined as the top 1% of attended patches, and HIPPO search algorithms were also used to identify the top 1% of patches by effect size.
@@
High attention regions drove counterintuitive effects in many specimens, while HIPPO-search-low-effect and HIPPO-search-high-effect identified more robust and consistent drivers of risk. High attention regions in high-risk cutaneous melanoma specimens (n=67) drove lower risk in 45% (n=30) of specimens. HIPPO-search-high-effect, on the other hand, identified regions that all drove higher risk and that more greatly contributed to high-risk predictions (, , independent t-test). High attention in high-risk breast cancer specimens (n=256) drove lower risk in 40% (n=102) specimens. Again, HIPPO-search-high-effect consistently identified regions that drove higher risk in the high-risk specimens (, , independent t-test) (Fig. 4a). High attention regions in low-risk SKCM specimens (n=67) drove higher risk in 10% (n=7). HIPPO-search-low-effect identified regions that all drove lower risk and more strongly contributed to lower risk predictions (, , independent t-test). High attention regions in low-risk BRCA specimens (n=256) drove higher risk predictions in 8% (n=20) specimens. HIPPO-search-low-effect identified patches that consistently drove lower risk predictions (, , independent t-test) (Fig. 4b). This counterintuitive effect underscores that attention scores may not directly relate to model predictions. Thus, interpretations that solely rely on these features may be misguided. HIPPO search algorithms reliably identified the regions that drove risk predictions and may have value as a tool for prognostic biomarker search.

### v3

@@ Results
We hypothesized that the low specificity of the UNI-based ABMIL model may be attributed to metastasis-induced alterations in the surrounding tumor microenvironment. To investigate this, we used HIPPO-knowledge to remove increasingly larger regions surrounding the tumor annotation and quantified the effect on metastasis detection. As the extent of peritumoral tissue removal increased, the UNI-based model was consistently more likely to predict the absence of metastasis (Supplementary Fig. LABEL:fig:supp_nec_dilations). Specificity increased from 0.73 at dilation of , to 0.78 at , to 0.80 at , to 0.86 at , and to 0.88 at . This was driven primarily by macrometastatic specimens, where specificity increased from 0.59 at dilation of to 0.68 at dilation of , to 0.73 at dilation of , to 0.82 at dilation of , to 0.86 at dilation of . Notably, other ABMIL models remained largely unaffected by peritumoral tissue removal, highlighting a unique characteristic of the UNI-based model. These observations may reflect a dependence of the UNI-based model on microenvironmental alterations associated with metastasis, though they may also indicate that the model has learned a less robust boundary for identifying metastatic disease. In summary, HIPPO enabled the quantitative exploration of peritumoral tissue on metastasis detection.
@@
While necessity assesses the importance of a feature or feature set, it does not inform whether the feature set is sufficient for model predictions. Metastasis detection models must be able to detect tumor regions no matter how small. Using the HIPPO-knowledge method with expert-annotated tumor regions as ground truth, we tested the sufficiency of metastatic regions using two methods: removing all non-tumor patches and measuring model outputs and adding tumor regions to normal specimens and measuring model outputs.
First, we constructed counterfactual specimens (n=49) by removing all non-tumor tissue (i.e., removing patches that did not intersect with expert tumor annotations) and measuring model outputs (Supplementary Fig. LABEL:fig:hippo-schematic-insert-deletea). Because these counterfactuals contained only tumor tissue, all images should be labeled “positive.” Under these conditions, the foundation models had the following sensitivity (true positive rate): UNI 0.98, REMEDIS 0.92, Phikon 0.98, CTransPath 0.96, RetCCL 0.82, Virchow2 0.96 (Fig. 2d).
By comparing these tumor-only predictions with those from the original WSIs, we assessed whether extratumoral tissue contributed to false negative errors. For micrometastatic specimens, five of the six foundation models showed improved sensitivity on the tumor-only images, suggesting that surrounding tissue influenced misclassification. Specifically, sensitivity increased by 25% for CTransPath, 4% for Phikon, 5% for REMEDIS, 100% for RetCCL, and 6% for Virchow2 . For UNI, the original WSIs yielded a perfect sensitivity for micrometastasis. However, when evaluated on tumor-only images, UNI generated a single false negative, suggesting that the model relied on extratumoral tissue to make a positive prediction. In this specimen, tumor tissue alone was insufficient for UNI to classify the slide as metastatic, indicating that surrounding regions were solely driving the positive prediction. A similar effect was observed with RetCCL in macrometastatic specimens: its true positive rate on the original WSIs was 0.95 (21 of 22 cases), whereas all macrometastases were correctly identified when evaluated on tumor-only images, demonstrating that extratumoral tissue caused the lone false negative in the original predictions.
@@
We also evaluated whether tumor itself was sufficient for metastasis detection by embedding tumor regions into normal specimens (Supplementary Fig. LABEL:fig:hippo-schematic-insert-deleteb). Specifically, we embedded all patches intersecting with tumor annotations into normal specimens, resulting in 3,920 positive counterfactual examples (80 normal slides 49 positive slides). Model outputs were then recorded for these tumor-in-normal counterfactuals (Fig. 2e). The resulting sensitivities were: UNI-based 0.98, REMEDIS-based 0.86, Phikon-based 0.95, CTransPath-based 0.90, RetCCL-based 0.63, and Virchow2-based 0.92 . Positive counterfactuals constructed with micrometastases were detected less reliably across most models (UNI-based achieved sensitivity of 0.96, REMEDIS-based 0.75, Phikon-based 0.91, CTransPath-based 0.93, RetCCL-based 0.40, Virchow2-based 0.85 ), suggesting that smaller tumors in the context of normal tissue are insufficient for positive metastasis detection.
@@
Finally, we quantified how strongly each tumor region generalizes across individuals by computing an average treatment effect for each metastatic slide. For each positive slide, we averaged the model’s predicted probability of metastasis across all tumor-in-normal counterfactuals, providing a measure of how reliably that tumor induces a positive prediction across diverse normal backgrounds. Using this metric, 100% of macrometastases (n=22) produced a positive average effect in the UNI-based, REMEDIS-based, Phikon-based, CTransPath-based, and Virchow2-based models. In the RetCCL-based model, 90% (20 of 22) of macrometastases produced a positive average effect. Micrometastases (n=27) were less likely to induce positive predictions on average, with 96% (n=26) positive in UNI, 93% (n=25) in Phikon, 81% (n=22) in CTransPath, 74% (n=20) in REMEDIS, 37% (n=10) in RetCCL, and 85% (n=23) in Virchow2 .
@@
To evaluate how metastasis size affects the sensitivity of ABMIL models, we analyzed metastasis-positive specimens from the CAMELYON16 test set. Our approach first involved initially removing all patch embeddings intersecting with expert tumor annotations, thereby creating metastasis-removed versions of each positive slide. effectively rendering the slide negative for metastases. We then introduced a single tumor region (shown in the right-hand side of Fig. 2a) into both 80 normal specimens and 49 metastasis-removed positive specimens. The REMEDIS-, Phikon-, and RetCCL-based ABMIL models detected 100% of counterfactuals as positive, demonstrating robustness to detecting even very small signature of tumor (Fig. 2f). In contrast, the UNI-based model failed to detect 41% (33 of 80) of positive counterfactuals, the CTransPath-based models failed 35% (28 of 80), and the Virchow2-based models failed 30% (24 of 80). A similar pattern was observed when the tumor region was embedded into the metastasis-removed positive specimens: REMEDIS-, Phikon-, and RetCCL-based models detected 100% of counterfactuals (49 of 49), while the UNI-based model missed 51% (25 of 49), the CTransPath-based model missed 65% (32 of 49), and the Virchow2-based model missed 37% (18 of 49) (Fig. 2f). This result is unexpected because the UNI-based model exhibited perfect sensitivity on the original CAMELYON16 test set (Fig. 2b) and achieved the highest sensitivity when larger tumors were embedded into normal tissue (Fig. 2e). This highlights that strong classification performance on the held-out test set does not necessarily reflect a model’s ability to generalize to more fine-grained or clinically relevant downstream applications.
@@
We also sought to quantify how sensitive each model is to individual tumor patches within positive specimens, as this can reveal whether different regions of tumor carry different levels of informativeness for machine learning classifiers. To accomplish this, we first removed all tumor patches intersecting with expert tumor annotations, producing a tumor-removed specimen. We then reintroduced patches that were fully contained within the expert tumor annotation, one at a time, into the tumor-removed specimen and evaluated the model outputs. These outputs were compared with the predictions made on the fully tumor-removed slide. While some individual tumor patches were sufficient to drive a positive prediction on their own, many were not (Fig. 2g for the UNI-based model, and other models and specimens are shown in Supplementary Figs. LABEL:fig:supp_all_single_patches_uni-LABEL:fig:supp_all_single_patches_virchow2).
@@
To further quantify how tumor size influences metastasis detection, we progressively added increasing numbers of tumor patches into normal slides and measured model sensitivity. All models demonstrated a graded response to tumor size, with sensitivity increasing as more tumor was added (Fig. 2h). UNI exhibited the highest sensitivity. For most models, sensitivity plateaued after approximately of tumor (16 patches) had been added. The RetCCL-based model showed the lowest overall sensitivity and was the least responsive to smaller tumor regions.
@@
To identify the largest amount of tumor that could be added while still going undetected, we applied a variant of HIPPO-search-negative-effect. Whereas the original HIPPO-search-negative-effect removes patches within a WSI, this variant progressively adds tumor patches to the “negative counterfactuals” described above. For each iteration, we selected the tumor patch whose effect size is lowest using the function. Using this procedure, we found that in some cases, regions up to could be added into a negative counterfactual without triggering a positive metastasis prediction. The tumor patches with the low effect sizes were largely similar in appearance to those that were sufficient to induce positive predictions, although some low-effect patches contained adipose cells along with tumor epithelial cells (Supplementary Fig. LABEL:fig:supp_mets_individual_tumor_patches_unseen). This shows that substantial regions of tumors can remain undetected by an ABMIL-based metastasis detection model. These biases warrant further investigation before these models are deployed in clinical settings.
@@
Given the influence of peritumoral tissue on UNI-based model predictions, we next evaluated whether peritumoral tissue were themselves sufficient to drive positive metastasis predictions. From each metastasis-positive specimen (n=49), we extracted halos of peritumoral tissue with widths of 64, 128, 256, or 1024 . Halos were defined beginning at either at the boundary of the expert tumor annotation or outside of the tumor annotation. Patches intersecting these halos were then added to normal specimens (n=80), resulting in 3,920 counterfactual examples (80 normal 49 positive specimens). Model predictions were averaged across the normal specimens to compute the average treatment effect for each peritumoral region. This was evaluated for the UNI-based and Phikon-based models.
@@
Halos of peritumoral tissue alone were sufficient to drive positive metastasis detection. Halos in width beginning at the tumor boundary caused positive predictions in 20% (n=10) of positive specimens and 10% (n=5) in the Phikon-based model (Supplementary Fig. LABEL:fig:supp_peritumoral). When the halo was shifted outside the tumor boundary, positive predictions occurred in 14% (n=7) of specimens for the UNI-based model and 10% (n=5) for the Phikon-based model. Even thin halos of beginning outside the tumor boundary triggered positive calls in 10% (n=5) of UNI-based cases and 6% (n=3) of Phikon-based cases (Supplementary Fig. LABEL:fig:supp_peritumoral). These results demonstrate that both models learned association between peritumoral tissue and metastasis status, despite the absence of any metastatic cells in these regions. Importantly, this bias could not have been uncovered using attention alone. HIPPO enabled the quantitative assessment of the influence of peritumoral tissue on metastasis detection across multiple foundation models.
HIPPO can help confirm regions that lead to false negative metastasis detections
@@
To further investigate patient samples that consistently produced false negative predictions, we visualized the attention maps of the ABMIL models to identify regions considered important by the model. In the CTransPath-based model, attention for specimen test_066 was primarily focused on adipose regions (Fig. 3a-b). Since attention maps only provide a qualitative visualization of regions in an image that the ABMIL models consider important, it is unclear to what extent adipose tissue directly affects model predictions. We address this with HIPPO-attention (Methods). The adipose regions were annotated and verified by an expert pathologist to ensure accurate identification. The adipose patches were then removed, and the effect of this perturbation on model outputs was quantified. Removing the adipose tissue restored the true positive prediction in specimen test_066, demonstrating that HIPPO can confirm which regions of a WSI are responsible for misclassification (Fig. 3c-3d).
To test whether the adipose tissue from test_066 could also induce misclassification in other models and specimens, we added the adipose regions from that specimen into each of the 48 remaining metastasis-positive specimens and recorded whether the added adipose caused false negatives (Supplementary Fig. LABEL:fig:hippo-schematic-insert-deleteb). We found that true positives were flipped to false negatives in 2, 2, 1, 1, and 5 specimens for UNI, REMIDIS, Phikon, CTransPath, and RetCCL, respectively (Fig. 3e-3f). This case study illustrates how HIPPO can elucidate biases that drive misclassification. While attention alone could not determine whether adipose tissue caused misclassification, it provided a useful hypothesis that we could formally evaluate with HIPPO-attention. This illustrates how HIPPO can complement attention-based interpretability analysis by quantitatively testing whether putatively important tissue regions truly influence model predictions.
@@
Identifying spurious correlations in deep learning models for medical imaging is essential for ensuring reliable and clinically meaningful predictions. To test HIPPO’s ability to identify spurious correlations, we deliberately introduced an artificial bias into the CAMELYON16 dataset (Supplementary Figs. LABEL:fig:cam16-shortcutsa and LABEL:fig:cam16-shortcutsb). Specifically, we inserted blue squares into all negative images, mimicking a plausible real-world scenario in which a pathologist marks certain slides with a blue marker. This modification creates a strong spurious correlation between the presence of a blue region and the negative label. We hypothesized that a model trained on this modified dataset would rely on the presence or absence of the blue region, rather than tumor morphology, to infer slide level labels.
@@
An ABMIL model using UNI embeddings was trained on the modified dataset and achieved a balanced accuracy of 1.0 on the test set, indicating that the spurious correlations turned the task into a trivial classification problem. By performing standard model interpretation using attention, we found that metastatic regions were considered highly important (Supplementary Fig. LABEL:fig:cam16-shortcutsc). However, removing these regions using HIPPO did not alter the model predictions, demonstrating that tumor regions were not important for model predictions despite a strong attention assignment. This highlights an important weakness of attention: the regions with high attention weights may not be the regions that actually drive predictions.
@@
Knowing that metastatic regions did not affect the model’s outputs, we used the search algorithm HIPPO-search-positive-effect (Methods) to identify the patches that maximally drove positive tumor predictions in both models for one positive specimen, test_002. Given that the model trained with spurious correlations uses the lack of a blue square as a cue for positive specimens, we expected that no individual patches would strongly drive the positive metastasis output and that tumor regions would not have a high effect on the prediction. Indeed, effect sizes were small and evenly distributed across the WSI (minimum , maximum , mean , and median ), indicating that no single region contributed strongly to the model prediction (Supplementary Fig. LABEL:fig:cam16-shortcutsb). By contrast, applying the same search algorithm to the model trained on the original CAMELYON16 datasetrevealed higher effect sizes (minimum , maximum , mean , and median ), with high-effect patches localized within expert tumor annotations (Supplementary Fig. LABEL:fig:cam16-shortcutse). We repeated this analysis across five additional positive specimens and consistently observed the same trends (Supplementary Table LABEL:table:supp_spurious).
These results show that simply inspecting performance and attention maps can obscure major failure modes. Although the model with the artificial shortcut achieved perfect accuracy and highlighted tumor regions in its attention map, these signals were misleading: tumor patches had almost no causal influence on the prediction. In contrast, HIPPO quantified the effect sizes of individual regions and correctly revealed that the model’s decisions were driven entirely by the introduced bias rather than by tumor morphology. By grounding interpretation in measured changes to model predictions, HIPPO provides a more faithful view of model behavior and can uncover hidden shortcut learning that conventional observational methods fail to detect.
@@
Having demonstrated HIPPO’s effectiveness in metastasis detection, where the regions of interest are well-defined and were previously annotated by expert pathologists, we extended our investigation to the more challenging task of cancer prognosis. Unlike metastasis detection, prognostic factors in WSIs are multifaceted and less clearly defined. We applied HIPPO to prognostic models that generate risk scores from WSIs, with the goal of identifying the tissue regions most responsible for these predictions.
@@
Our experiments yielded two key insights. First, HIPPO’s search algorithms demonstrated superior ability in identifying tissue patches that consistently and significantly influence risk predictions compared to conventional attention. Attention-based interpretations often produced mixed effects or counterintuitive effects, such as highlighting regions that lowered predicted risk in otherwise clearly high-risk specimens. HIPPO, on the other hand, provided a consistent, reliable, and quantitative assessment of the regions that drive risk. Second, HIPPO’s perturbation-based design enabled “virtual experiments” to assess the effect of targeted tissue interventions on prognostic outcomes in silico through the lens of the ABMIL model. This capability positions HIPPO as a promising tool for accelerating the discovery and validation of prognostic tissue biomarkers, thereby narrowing the gap between computational analysis and clinical actionability.
@@
We trained prognostic ABMIL models using the PORPOISE framework [chen2022porpoise], a computational tool designed for predicting survival outcomes from histopathology images, to predict overall survival from WSIs in breast cancer (TCGA-BRCA) and cutaneous melanoma (TCGA-SKCM). The same training and validation splits were used as in the original publication. Non-overlapping patches from WSIs were embedded using the UNI model [chen2024uni] (in the original PORPOISE publication, a truncated ResNet50 [he2016resnet] was used). The resulting models successfully stratified patients into low- and high-risk groups, achieving concordance indices of 0.667 and 0.557 for TCGA-BRCA and TCGA-SKCM, respectively (Supplementary Fig. LABEL:fig:supp_surv_km). Low and high risk were defined as the first and fourth quartiles of risk scores. High attention regions were defined as the top 1% of attended patches. For comparison, HIPPO search algorithms was used to identify the top 1% of patches based on effect size.
@@
High attention regions drove counterintuitive effects in many specimens, while HIPPO-search-positive-effect and HIPPO-search-negative-effect identified more robust and consistent drivers of risk. High attention regions in high-risk cutaneous melanoma specimens (n=67) drove lower risk in 45% (n=30) of specimens. HIPPO-search-positive-effect, on the other hand, identified regions that all drove higher risk and that more greatly contributed to high-risk predictions (, , independent two-sided t-test). High attention in high-risk breast cancer specimens (n=256) drove lower risk in 40% (n=102) specimens. Again, HIPPO-search-positive-effect consistently identified regions that drove higher risk in the high-risk specimens (, , independent two-sided t-test) (Fig. 4a). High attention regions in low-risk SKCM specimens (n=67) drove higher risk in 10% (n=7). HIPPO-search-negative-effect identified regions that all drove lower risk and more strongly contributed to lower risk predictions (, , independent two-sided t-test). High attention regions in low-risk BRCA specimens (n=256) drove higher risk predictions in 8% (n=20) specimens. HIPPO-search-negative-effect identified patches that consistently drove lower risk predictions (, , independent two-sided t-test) (Fig. 4b). These counterintuitive effects underscore that attention scores do not necessarily correspond to a region’s influence on the model output, and that interpretations relying solely on attention may be misguided. In contrast, HIPPO search algorithms robustly identified the regions that drove risk predictions and may be valuable tools for discovering prognostic biomarker.